OpenStudy (anonymous):

Factor theorem. what is the polynomial function in standard form with zeros? -1,2,3 0,1,-2,-3

OpenStudy (anonymous):

for the first one it is the one that looks like \[(x+1)(x-2)(x-3)\] after you multiply out

OpenStudy (anonymous):

if you do it carefully you should get \[x^3-4 x^2+x+6\]

OpenStudy (anonymous):

second one is similar, but this time you have to multiply \[x(x-1)(x+2)(x+3)\]

OpenStudy (anonymous):

how did u get the first answer can u explain or can it be explained?

OpenStudy (anonymous):

you have to multiply it out, one at a time, using the distributive law or perhaps something you heard called "foil"

OpenStudy (anonymous):

\[(x+1)(x-2)=x^2-2x+x-2=x^2-x-2\] then \[(x^2-x-2)(x-3)=x^3-x^2-2x-3x^2+3x-6\]\[=x^3-4x+x-6\]

OpenStudy (anonymous):

ok so what about the second one since you said i have to multiply